Stark is seeking to hold his $20,000 savings in a savings account for 1 year. He enquires with three banks, Bank A, Bank B, and Bank C. Bank A offers J12 = 12%; Bank B offers J2 =12.1%, and Bank C offers J4 = 12.05%. Which bank should Stark deposit his savings?
1
Expert's answer
2020-06-17T19:29:15-0400

A

"FV=20 000(1+\frac{12}{12})^{12}=22 536.50"


B

"FV=20 000(1+\frac{12.1}{2})^2=22 493.21"


C


"FV=20 000(1+\frac{12.05}{4})^4= 22521.11"


 Stark deposit his savings to bank A.